OpenStudy (anonymous):

determine the 10th term in the pattern : 5 , 8 11 ,14 , 17....

OpenStudy (anonymous):

do you see the difference between the terms?

OpenStudy (anonymous):

@Silver*Dancer, do you see the difference between the terms, if no I can explain what it is.....

OpenStudy (anonymous):

It is an A.P tn=a+(n-1)d a=5,n=10, d=8-5=3 calculate t10

OpenStudy (anonymous):